The sum of the distances of a moving point from (3, 0) and (-3, 0) is 8. Find its equation


1
Expert's answer
2021-10-15T10:00:07-0400

Let the coordinates of the moving point be "(x,y)."

Therefore, according to the question:

"\sqrt{(x-3)^2+y^2}+\sqrt{(x+3)^2+y^2}=8\\
\Rightarrow \sqrt{(x-3)^2+y^2}=8-\sqrt{(x+3)^2+y^2}"

Squaring both sides, we get:

"(x-3)^2+y^2=64+(x+3)^2+y^2-16\sqrt{(x+3)^2+y^2}\\
\Rightarrow -6x=64+6x-16\sqrt{(x+3)^2+y^2}\\
\Rightarrow 16\sqrt{(x+3)^2+y^2}=12x+64\\
\Rightarrow 4\sqrt{(x+3)^2+y^2}=3x+16"

Squaring both sides, we get:

"16(x+3)^2+16y^2=9x^2+256+96x\\
\Rightarrow 7x^2+16y^2=112\\
\Rightarrow \frac{x^2}{16}+\frac{y^2}{7}=1"